how many pounds of a 35% salt solution and how many pounds of 14% salt solution should be combined so that 50 pounds of a 20% salt solution will be obtained?
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6

Answer by ewatrrr(24785)

.35x + .14(50lb-x) = .20(50lb)
x = .06(50lb)/.21, amount of 35% salt solution
50-x = amount of 14% salt solution